OpenStudy (isaacroberts242):

Find f(-11/5) if f(n) = 5n + 6.

OpenStudy (anonymous):

substitute -11/5 for n \[f \left( \frac{ -11 }{ 5 } \right)=5\left( \frac{ -11 }{ 5 } \right)+6\]

OpenStudy (anonymous):

cancel the 5's, then add